### v1

# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) on weights and discuss SORSA’s superiority in minimizing the deviation from the pre-trained weight. Each SORSA layer consists of two main parts: trainable principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement an orthonormal regularizer and analyze its importance by performing gradient analysis. The analysis shows that the regularizer could effectively transfer the scaling information into $\Sigma_{p}$, which ensures the parameter updating of SORSA layers is evenly and minimized on $U_{p}$ and $V^{\top}_{p}$. SORSA layers could be merged during inference, thus eliminating inference latency. After all, SORSA shows a faster convergence speed than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%) We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
Deep Learning, LLM, PEFT, LoRA, SVD
Figure 1: Architecture of a SORSA layer. We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
1 Introduction
Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks. For specific downstream tasks, full parameter fine-tuning, which continues training all parameters of LLMs on downstream data, is widely used.
However, as the number of parameters in LLMs rapidly increases, full parameter fine-tuning becomes increasingly inefficient. For example, the estimated VRAM requirement for fully fine-tuning Llama 2 7B using Float32 could approach approximately 100 GB, making it unlikely to fully fine-tune the model on a single GPU with current technology. Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB, thus rendering it unfeasible on a single GPU with current technology.
To address these challenges, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021) have been proposed. These methods enable the training of only a limited number of parameters, which significantly reduces VRAM requirements while achieving comparable or even superior performance to full fine-tuning. For instance, tuning Llama 2 7B in Float32 by LoRA with a rank of 128 only takes approximately 60GB VRAM, which allows training in 1 $\times$ NVIDIA A100-80GB or 3 $\times$ NVIDIA RTX 4090-24GB.
Among those PEFT methods, LoRA (Hu et al., 2021) and its variants (Zhang et al., 2023; Meng et al., 2024; Liu et al., 2024; Dettmers et al., 2024) had become increasingly popular due to their: 1. low training VRAM requirement 2. no inference latency 3. versatility in different neuron network architectures.
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA layer has two main parts: principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principle singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA layer is illustrated in Figure 1.
Furthermore, we analyze the pattern of variance of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
In our experiments, SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning.
2 Related Works
Parameter-efficient fine-tuning (PEFT) methods have been developed to address the inefficiency of full parameter fine-tuning for large language models. These methods focus on adapting the model for downstream tasks while updating only a few parameters and keeping most of the model’s weights frozen. This approach significantly reduces the memory and computational requirements during training, especially VRAM.
Adapter-based PEFT is the first type of PEFT, which was initially designed by (Houlsby et al., 2019). It introduces additional trainable non-linear blocks into the frozen pre-trained model, which could effectively tune the pre-trained model with a limited amount of trainable parameters. Its variants, e.g., (Lin et al., 2020), reduce the number of adapter layers per block, and (He et al., 2022) focus on adding adapter modules parallel to existing layers. However, all adapter-based PEFT methods introduce inference latency due to their non-mergeable attribute.
Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in (Lester et al., 2021). There are several variants of this work, including (Liu et al., 2022; Razdaibiedina et al., 2023). However, they have some inevitable shortcomings, such as potential performance limitations compared to full parameter fine-tuned models, additional inference latency due to expanding the length of the total input to the model, and the complexity of designing effective initialization.
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating; DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA; PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}$ into frozen $W_{pri}=AB$ and $W_{res}=W_{0}-W_{pri}$ based on singular values, and only updating the most important part $W_{pri}$, which result a faster convergence speed and better fitting than LoRA. SORSA has a similar architecture to PiSSA, which both conducts singular value decomposition (SVD) and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
Other methods. There are also a few PEFT methods with unique techniques. For example, GaLore (Zhao et al., 2024) is a memory-efficient PEFT method that reduces VRAM usage by leveraging gradient accumulation and low-rank approximation. Despite its efficiency, GaLore’s complex implementation, involving Singular Value Decomposition (SVD) on gradients, adds computational complexity and may face scalability issues for very large models. LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance. However, LISA’s selective fine-tuning requires careful resource allocation and introduces additional training complexity, complicating the pipeline and necessitating more sophisticated monitoring mechanisms.
3 Singular Values and Vectors Analysis
3.1 Singular Value Decomposition
The geometric meaning of SVD can be summarized as follows: for every linear mapping $W:K^{m}\to K^{n}$, the singular value decomposition finds a set of orthonormal bases in both the original space and the image space such that $W$ maps the $i$-th basis vector of $K^{m}$ to a non-negative multiple of the $i$-th basis vector of $K^{n}$, and maps the remaining basis vectors of $K^{m}$ to the zero vector. In other words, the matrix $W$ can be represented as $\Sigma$ on these selected bases, where $\Sigma$ is a diagonal matrix with all non-negative diagonal elements.
According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\Sigma V^{\top}$. Here, $U\in\mathbb{R}^{m\times k}$ are left singular vectors and have orthonormal columns, $V^{\top}\in\mathbb{R}^{k\times n}$ are right singular vectors and have orthonormal rows, and $\Sigma\in\mathbb{R}^{k\times k}$ is a diagonal matrix, where diagonal values are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order.
3.2 Analysis Method
The study of DoRA (Liu et al., 2024) introduces an analysis method that focuses on the deviation of magnitude and direction ($\Delta M,\Delta D$) during training of full parameter fine-tuning and LoRA (Hu et al., 2021). They discovered that the distinction between full parameter fine-tuning and LoRA likely affects their learning ability difference. Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices. Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m,n}$ and tuned weight $W_{t}\in\mathbb{R}^{m,n}$, which $t$ denotes the training step, could be defined as follows:
$\Delta\Sigma_{t}=\frac{\Sigma_{i=1}^{k}\left|\sigma_{t}^{i}-\sigma_{0}^{i}\right|}{k}$ (1)
Here, $\Delta\Sigma_{t}$ represents singular value variants between $W_{0}$ and $W_{t}$ at training step $t$. $\sigma^{i}_{t}$ denotes the $i$-th element in diagonal of $\Sigma_{t}$, where $\Sigma_{t}$ is decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$;
$\Delta U_{j}=\left|\sum_{i=1}^{m}\left(U_{t}\odot U_{0}\right)_{ij}\right|$ (2)
$\Delta V^{\top}_{i}=\left|\sum_{j=1}^{n}\left(V_{t}^{\top}\odot V_{0}^{\top}\right)_{ij}\right|$ (3)
$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{i}+\Delta V^{\top}_{i})$ (4)
Here, $\Delta D_{t}\in(0,1)$ represents variants of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$. $\Delta U\in\mathbb{R}^{k}$, $\Delta V^{\top}\in\mathbb{R}^{k}$. $\odot$ denotes element-wise product, and $\left|\cdot\right|$ denotes element-wise absolute value. $U_{t}$ and $V_{t}$ are decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$.
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer with $\gamma=0.1$ and without regularizer). See Appendix A.1 for training details of the analysis.
3.3 Analysis Result
This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect. The analysis data is illustrated in Figure 2. The four methods obtained similar final loss ($\pm 10\%$).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: $\Delta D$ and $\Delta\Sigma$ of each trainable parameters during training steps. Numbers in the plot represent layer of the weight. Dots represent mean $\Delta D$ and $\Delta\Sigma$ at specific step.
The partial fine-tuning and LoRA methods analysis show that both methods exhibit significant adjustments in significant vectors $\Delta D$. This substantial alteration disrupts the characteristics of the pre-trained matrix and is likely to affect the model’s generalization ability. Moreover, the updates of all parameters in both partial fine-tuning and LoRA methods show a parallel updating pattern across weights in different layers, which emphasizes a restriction of these methods and may lead to a potentially disruptive modification of the generalization ability.
SORSA with regularizer shows $\Delta D$ and $\Delta\Sigma$ are in a much smaller range than other analysis methods. Moreover, different matrices demonstrated unrelated updating patterns, compared to all parallel-like other three methods, indicating that the updates in the SORSA are less constrained and potentially converge faster. In contrast, SORSA without the orthonormal regularizer exhibits a greater $\Delta D$ and $\Delta\Sigma$ alteration at each training step compared to SORSA with the regularizer. SORSA without regularizer also shows a linear-like updating pattern, similar to LoRA and partial FT. The result verifies the importance of regularizer in maintaining stability and minimizing the deviation from pre-trained matrices during the training process. These updating patterns indicate that SORSA maintains the characteristics of the pre-trained matrix better, thus potentially preserving the model’s generalization ability more effectively. This property allows SORSA layers to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
4 Our Method
4.1 Singular Value and Orthonormal Regularized Singular Vector Adaptation
According to the definition of SVD in Section 3.1, given a rank $r$ where $r\ll k$, we could perform the low-rank approximation by selecting the first $r$ items on the diagonal of $\Sigma$, which is the first $r$ most significant singular values, and also select the first $r$ columns of $U$ and first $r$ rows of $V^{\top}$, which correspond to the selected singular values. By performing SVD low-rank approximation, we could get a low-rank matrix that preserves the largest significant values and vectors, containing the matrix’s most significant data.
Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principle weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
$W_{0}=W_{p}+W_{r}$ (5)
$W_{p}=U_{[:,:r]}\Sigma_{[:r,:r]}V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (6)
$W_{r}=U_{[:,r:]}\Sigma_{[r:,r:]}V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (7)
Here, $U$ represents the left singular vectors, $\Sigma$ represents the diagonal matrix of singular values, and $V^{\top}$ represents the right singular vectors. We use PyTorch (Paszke et al., 2019) syntax to demonstrate matrix selection, where $[:,:r]$ denotes selecting the first $r$ columns of the matrix, and $[r:,:]$ denotes selecting the last $r$ rows of the matrix. We rewrite $U_{[:,:r]}$, $\Sigma_{[:r,:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $\Sigma_{[r:,r:]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$ correspondingly.
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Eq (8), which is initially equivalent to the pre-trained weight $W_{0}$.
During training, $W_{r}$ remains frozen, and only $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ are updated. Since $\Sigma_{p}$ is a diagonal matrix, only the diagonal elements are updated, while the rest of the matrix remains zero.
$\text{SORSA}(x):=W_{r}x+W_{p}x$ (8)
We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 4.2.
$\mathcal{L}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V_{p}V^{\top}_{p}-I\|_{F}$ (9)
Where $\mathcal{L}_{reg}$ is the orthonormal regularizer loss, the $U_{p}$ and $V^{\top}_{p}$ are each orthonormal vectors in columns and rows, respectively, after initialization due to SVD’s property. The regularizer could keep their orthonormality during training.
Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA layer at training step $t$ could be expressed as:
$W_{p,t+1}=$ $W_{p,t}-\eta_{t}\nabla_{W_{p,t}}\mathcal{L}_{train}-\gamma_{t}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (10)
At training step $t$, $\nabla_{W_{t}}\mathcal{L}_{train}$ denotes the gradient of $W_{t}$ respect to $\mathcal{L}_{train}$, and $\nabla_{W_{t}}\mathcal{L}_{reg}$ denotes the gradient of $W_{t}$ respect to the orthonormal regularizer loss $\mathcal{L}_{reg}$. $\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss, respectively.
We update the SORSA as Eq (11) for implementation simplicity.
$W_{p,t+1}=$ $W_{p,t}-\eta_{t}\left(\nabla_{W_{p,t}}\mathcal{L}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}\mathcal{L}_{reg}\right)$ (11)
$\eta_{d}$ is the learning rate which is given to the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
4.2 Gradient Analysis
In this section, we analyze the impact of an orthonormal regularization on the updates of the singular vector matrix $U_{p}$ and $V^{\top}_{p}$ during training and the importance of keeping $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ separately. We also analyze the scaling factor for the learning rate of $\Sigma_{p}$
Given $U_{p,t-1}\in\mathbb{R}^{m\times r}$ is an orthonormal matrix, which corresponds to the initial state of $U_{p}$ due to the property of SVD. Here, we denote the gradient without regularizer for matrix $W$ at step $t$ is by $\nabla_{W_{t}}$, and the gradient with regularizer by $\nabla^{\prime}_{W_{t}}$:
$\nabla_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (12)
$\nabla^{\prime}_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}+\gamma\nabla_{U_{p,t-1}}\mathcal{L}_{reg}$ (13)
Where $\nabla_{U_{p,t-1}}\mathcal{L}_{reg}=\frac{\partial\mathcal{L}_{reg}}{\partial U_{p,t-1}}=0$ due to the orthonormality of $U_{p,t-1}$; $\nabla_{U_{p,t-1}}\mathcal{L}_{train}=\frac{\partial\mathcal{L}_{train}}{\partial U_{p,t-1}}$.
Thus, we could calculate the weight at step $t$ without regularizer $U_{t}$, and with regularizer $U_{t}^{\prime}$:
$U_{p,t}^{\prime}=U_{p,t}=U_{p,t-1}-\eta\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (14)
Based on their weight on step $t$, we could calculate $W_{t+1}$ and $W^{\prime}_{t+1}$. Here, we only focus on the gradient of $U_{p}$ and $V^{\top}_{p}$, thus ignoring the updating of $\Sigma_{p,t}$ in this step.
$U_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}$ (15)
$V^{\top}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}$ (16)
$W_{t+1}=U_{t+1}\Sigma_{p,t+1}V^{\top}_{t+1}$ (17)
$U^{\prime}_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}$ (18)
$V^{\top\prime}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}$ (19)
$W^{\prime}_{t+1}=U^{\prime}_{t+1}\Sigma_{p,t+1}V^{\top\prime}_{t+1}$ (20)
Under optimal circumstances, $U^{\prime}_{p,t+1}$ will be an orthonormal matrix. Therefore, $U^{\prime}_{p,t+1}$ can be interpreted as pure rotation, which does not change the norm of the input matrix. By enforcing the orthonormal property of $U_{p}$ and $V^{\top}_{p}$ at each step, their regularized gradients $\nabla_{U_{p}}\mathcal{L}_{reg}$ and $\nabla_{V^{\top}_{p}}\mathcal{L}_{reg}$ will ensuring that the scaling information is not propagated into $U_{p,t}$ and $V^{\top}_{p,t}$. Consequently, only $\Sigma_{p}$ retains all scaling information.
To prove this statement, we could calculate the difference between $W^{\prime}_{t+1}$ and $W_{t+1}$:
$\begin{split}W^{\prime}_{t+1}-W_{t+1}&=-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}U_{p,t}\Sigma_{p,t+1}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{train}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\\ &\quad-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}V^{\top}_{p,t}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}\Sigma_{p,t+1}\\ &\quad+\gamma^{2}\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\end{split}$ (21)
Here, we could find that each term has the $\Sigma_{p,t+1}$. This shows that by optimizing respected to regularizer loss at step $t+1$ and optimizing $\Sigma_{p,t+1}$ respected to train loss on step $t+2$, we could efficiently increase the orthonormality of $U_{p}$ and $V^{\top}_{p}$ while have negligible side-effect on convergence. In short, adding regularizer could transfer scaling information from $U_{p}$ and $V^{\top}_{p}$, which come from optimizing for $\mathcal{L}_{train}$, into $\Sigma_{p}$ allows the matrix scaling information updates more holistically, ultimately presenting a better generalization. In Figure 3, we illustrate the effect of $\gamma$ on orthonormality, training loss, and gradient norm. The figures present almost no difference in loss and gradient norm in different $\gamma$ values, which indicates the strong robustness of training of SORSA.
Figure 3: Orthonormality, loss and gradient norm variation of SORSA using different $\gamma$ on MetaMathQA training of Mistral 7B v0.1.
Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$. By ensuring the scaling information only be updated into $\Sigma_{p}$, we could also ensure that the parameter distribution of $W_{p}$ is evenly distributed.
Introducing the orthonormal regularizer into the training process promotes a more stable update trajectory and a more evenly distributed weight by actively countering deviations from orthonormality. This regulation stabilizes the updates and enhances the model’s ability to maintain crucial properties of the singular vectors $U_{p}$ and $V^{\top}_{p}$, thereby contributing to improved generalization and effectiveness in learning. This analysis underlines the utility of orthonormal regularization in complex model training scenarios, particularly in fine-tuning pre-trained models where preserving learned features is crucial.
5 Empirical Experiments
We conduct comparative experiments on different NLP tasks, including natural language generation (NLG) between SORSA, PiSSA (Meng et al., 2024), LoRA (Hu et al., 2021), and full parameter fine-tuning.
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023). We trained the models with SORSA applied and conducted GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) tests. The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). See Appendix A.2 for more details and hyperparameters of the training. We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from the PiSSA paper (Meng et al., 2024). For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80G) GPU. See Table 1 for the results and Figure 4 for the loss and gradient norm comparison. See Figure 5 for the variation in orthonormality of SORSA.
Model Method Trainable Parameters MATH GSM-8K
Llama 2 7B Full FT 6738M 7.22 49.05
LoRA 320M 5.50 42.30
PiSSA 320M 7.44 53.07
SORSA 320M 10.36 56.03
Mistral 7B Full FT 7242M 18.60 67.02
LoRA 168M 19.68 67.70
PiSSA 168M 21.54 72.86
SORSA 168M 21.86 73.09
Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from the paper of PiSSA (Meng et al., 2024).
Refer to caption Refer to caption
Figure 4: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B.
Figure 5: Variation of orthonormality on MetaMathQA training of Llama 2 7B and Mistral 7B v0.1.
The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT. On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both tasks compared to the other three methods.
In this section, due to the limitation of computing resources, we train and benchmark a small number of tasks.
6 Conclusion
In this paper, we introduced SORSA, a novel parameter-efficient fine-tuning (PEFT) method designed to enhance the adaptation of large language models (LLMs) for downstream tasks. SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principle and residual components, only training the principle singular values and vectors while freezing the residuals. We implemented an orthonormal regularizer to maintain the orthonormality of singular vectors during training, ensuring efficient parameter updates and preserving the integrity of singular values.
Our experiments demonstrated that SORSA outperforms existing PEFT methods, such as LoRA and PiSSA, in both convergence speed and accuracy on the NLG tasks. Specifically, Llama 2 7B, tuned with SORSA, achieved significant improvements in the GSM-8K and MATH benchmarks, highlighting the effectiveness of our approach.
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis. We also introduce the gradient analysis of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
SORSA retains the advantages of LoRA and variants, including low training VRAM requirements, no inference latency, and versatility across different neural network architectures. By offering a more efficient fine-tuning mechanism, SORSA presents a promising direction for future research and application in the field of LLMs.
Overall, SORSA gives a new perspective on parameter-efficient fine-tuning, showcasing exceptional efficiency and robust performance. It not only outperforms existing methods like LoRA and PiSSA in several downstream tasks but also maintains the practical benefits of low VRAM requirements, no inference latency, and ease of implementation. This innovative approach offers a promising direction for future research and practical applications in adapting pre-trained models, making it a pivotal development in the field.
7 Future Work
While SORSA demonstrates substantial improvements over existing PEFT methods, there are several avenues for future work to enhance its capabilities further and extend its applicability:
  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different transformer-based LLMs. This includes replicating other tests from studies like PiSSA (Meng et al., 2024) and performing evaluations such as HumanEval (Chen et al., 2021) and MT-Bench (Zheng et al., 2023), in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2021a) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
  • Application to Other Fields: Test SORSA on different fields beyond NLP. Such as conducting experiments on computer vision (CV) models like DDPM (Ho et al., 2020), DDIM (Song et al., 2020), and ViT (Dosovitskiy et al., 2021). These experiments could expand SORSA’s applicability to a broader range of deep-learning tasks.
  • Quantized SORSA (QSORSA): Investigate the application of quantization techniques to SORSA introduced in QLoRA (Dettmers et al., 2024). This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, QSORSA could enable highly efficient and powerful language models suitable for a broader range of practical applications.
By pursuing these directions, future work can build on the foundation laid by SORSA, pushing the boundaries of PEFT and enhancing the adaptability and performance of large language models across a wide range of applications. This will enable the development of more versatile downstream models, potentially expanding the impact of Machine Learning into more fields and becoming an integral part of people’s everyday lives.
8 Impact Statement
In this paper, we introduce an innovative PEFT method in the area of Machine Learning. Our approach significantly streamlines the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability. As we push the boundaries of what is possible with machine learning, it is essential to consider the broader impacts of these advancements on both the environment and ethical standards within the field.
8.1 Environmental Impact
SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B model than full parameter fine-tuning. This significant reduction in hardware resource requirements translates to approximately 50% less energy consumption compared to full parameter fine-tuning methods. By enhancing efficiency, our approach contributes to reducing the environmental footprint of machine learning operations.
8.2 Ethical Concerns
The PEFT method, while efficient, raises critical ethical concerns regarding the security of built-in safety measures in AI models. As demonstrated in (Lermen & Rogers-Smith, 2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content. The ease and affordability of such methods underscore the vulnerability of safety protocols. It is imperative to develop robust safeguards that keep pace with technological advancements, ensuring that efficiency gains in model tuning do not compromise the ethical use of AI.
9 Acknowledgment
We would like to extend our gratitude to Prof. Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing guidance to the research.
Appendix A Training Details
A.1 Analysis
For the singular values and vectors analysis in section 3, we apply partial fine-tuning, LoRA SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, using the same training data and optimizer as A.2. In the analysis, LoRA and SORSA are only applied to q_proj and v_proj matrices, respectively. FT is only trainable on q_proj and v_proj. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 2 for hyperparameters.
Training
Method FT LoRA SORSA (w/o reg) SORSA
Precision TF32 + BF16 Mix-Precision
Epoch 1
Batch Size 128
Max Length 512
Weight Decay 0.00
Warm-up Ratio 0.03
Learning Rate 2e-5 2e-5 2e-5 3e-5
$\gamma$ 0 5e-4
Rank 128 128 128
Table 2: Hyperparameters for the analysis
A.2 NLG Experiments
For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA layers are applied on all linear matrices in every layer. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 3 for hyperparameters.
Training
Model Llama 2 7B Mistral 7B v0.1
Method SORSA
Precision TF32 + BF16 Mix-Precision
Epoch 1
Batch Size 128
Max Length 512
Weight Decay 0.00
Warm-up Ratio 0.03
Learning Rate 3e-5
$\gamma$ 4e-4
Rank 128 64
Inference
Precision BF16
Top-P 1.0
GSM-8K Max Length 1024
MATH Max Length 2048
Table 3: Hyperparameters of training on models for GSM-8K and MATH