### v4

# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect. Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\text{diag}(S_{p})V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\text{diag}(S_{r})V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement and analyze an orthonormal regularizer, which we prove could decrease the condition number of $W_{p}$ and allows the optimization to be more efficient. SORSA adapters could be merged during inference, thus eliminating any inference latency. After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments. On the GSM-8K benchmark, Llama 2 7B adapted using SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%). On the MATH benchmark, SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
Figure 1: Architecture of a SORSA adapter. We only train parts rendered in orange ($U_{p}$, $\text{diag}(S_{p})$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\text{diag}(S_{r})$ and $V^{\top}_{r}$).
1 Introduction
Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks (Peng et al., 2024; Touvron et al., 2023; Dubey et al., 2024; Radford et al., 2019; OpenAI, 2023). For specific downstream tasks, full parameter fine-tuning, which continues training all parameters of LLMs on downstream data, is widely used.
However, as the number of parameters in LLMs rapidly increases, full parameter fine-tuning becomes increasingly inefficient. For example, the estimated VRAM requirement for fully fine-tuning Llama 2 7B using Float32 could approach approximately 100 GB, making it unlikely to fully fine-tune the model on a single GPU with current technology. Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB (Touvron et al., 2023; Anthony et al., 2023), thus rendering it unfeasible on a single GPU with current technology.
To address these challenges, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021) have been proposed. These methods enable the training of only a limited number of parameters, which significantly reduces VRAM requirements while achieving comparable or even superior performance to full fine-tuning. For instance, tuning Llama 2 7B in Float32 by LoRA (Hu et al., 2021) with a rank of 128 only takes approximately 60GB VRAM, which allows training on 1 $\times$ NVIDIA A100 (80GB), or even 3 $\times$ NVIDIA RTX 4090 (24GB).
Among those PEFT methods, LoRA (Hu et al., 2021) and its variants (Zhang et al., 2023; Meng et al., 2024; Liu et al., 2024; Dettmers et al., 2024) had become increasingly popular due to their: 1. Low training VRAM requirement 2. No inference latency 3. Versatility in different neuron network architectures.
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\text{diag}(S_{p})V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\text{diag}(S_{r})V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principal singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA adapter is illustrated in Figure 1.
Furthermore, we analyze the pattern of variation of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
We also provide a comprehensive gradient analysis that provides a mathematical foundation for SORSA. This analysis demonstrates several crucial properties of our method, including the convexity of the regularizer, Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$. Moreover, we prove that SORSA improves the condition number of the optimization problem compared to unregularized approaches.
SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning in our experiments.
2 Related Works
Parameter-efficient fine-tuning (PEFT) methods have been developed to address the inefficiency of full parameter fine-tuning for large language models. These methods focus on adapting the model for downstream tasks while updating only a few parameters and keeping most of the model’s weights frozen. This approach significantly reduces the memory and computational requirements during training, especially VRAM.
Adapter-based PEFT is the first type of PEFT, which was initially designed by Houlsby et al. (2019). It introduces additional trainable non-linear blocks into the frozen pre-trained model, which could effectively tune the pre-trained model with a limited amount of trainable parameters. Its variants, e.g., Lin et al. (2020), reduce the number of adapter layers per block, and He et al. (2022) focus on adding adapter modules parallel to existing layers. However, all adapter-based PEFT methods introduce inference latency due to their non-mergeable attribute.
Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in Lester et al. (2021). There are several variants of this work, including Liu et al. (2022); Razdaibiedina et al. (2023). However, they have some inevitable shortcomings, such as potential performance limitations compared to full parameter fine-tuned models, additional inference latency due to expanding the length of the total input to the model, and the complexity of designing effective initialization.
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating. DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA. OLoRA (Büyükakyüz, 2024) uses QR decomposition to initialize the LoRA adapters $A$ and $B$, which initializes $B$ as an orthogonal matrix. They discusses the significance of orthonormality in neural networks’ weight (See Section 5 for more details). In their experiments, OLoRA demonstrates faster convergence than LoRA. PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=AB$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen. PiSSA results a faster convergence speed and better fitting compared to LoRA. SORSA has a similar architecture to PiSSA, which both conducts SVD and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
Other methods. There are also a few efficient adapting methods with unique techniques. For example, GaLore (Zhao et al., 2024) is a memory-efficient PEFT method that reduces VRAM usage by leveraging gradient accumulation and low-rank approximation. Despite its efficiency, GaLore’s complex implementation, involving perform SVD on gradients, adds computational complexity and may face scalability issues for very large models. LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance, and selectively fine-tune essential parameters.
3 Singular Values and Vectors Analysis
3.1 Singular Value Decomposition
The geometric meaning of SVD can be summarized as follows: for every linear mapping $W:K^{m}\to K^{n}$, the singular value decomposition finds a set of orthonormal bases in both the original space and the image space such that $W$ maps the $i$-th basis vector of $K^{m}$ to a non-negative multiple of the $i$-th basis vector of $K^{n}$, and maps the remaining basis vectors of $K^{m}$ to the zero vector. In other words, the matrix $W$ can be represented as $\Sigma$ on these selected bases, where $\Sigma$ is a diagonal matrix with all non-negative diagonal elements.
According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\text{diag}(S)V^{\top}$. Here, $U\in\mathbb{R}^{m\times k}$ is matrix of left singular vectors and have orthonormal columns, $V\in\mathbb{R}^{n\times k}$ is matrix of right singular vectors and have orthonormal columns, and $S\in\mathbb{R}^{k}$ are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order. $\text{diag}(W)$ denotes a function to form a diagonal matrix from $W$.
3.2 Analysis Method
The study of DoRA (Liu et al., 2024) introduces an analysis method that focuses on the deviation of magnitude and direction ($\Delta M,\Delta D$) during training of full parameter fine-tuning and LoRA (Hu et al., 2021). They discovered that the distinction between full parameter fine-tuning and LoRA likely affects their learning ability difference. Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices during updating. Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$ and tuned weight $W_{t}\in\mathbb{R}^{m\times n}$, which $t$ denotes the training step, could be defined as follows:
$\Delta\Sigma_{t}=\frac{\Sigma_{i=1}^{k}\left|\sigma_{t}^{i}-\sigma_{0}^{i}\right|}{k}$ (1)
Here, $\Delta\Sigma_{t}$ represents singular value variants between $W_{0}$ and $W_{t}$ at training step $t$. $\sigma^{i}_{t}$ denotes the $i$-th element in diagonal of $\Sigma_{t}$, where $\Sigma_{t}$ is decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$;
$\Delta U_{t,j}=\left|\langle{\mathbf{u}}_{t}^{j},{\mathbf{u}}_{0}^{j}\rangle\right|$ (2)
$\Delta V^{\top}_{t,i}=\left|\langle{\mathbf{v}}_{t}^{i},{\mathbf{v}}_{0}^{i}\rangle\right|$ (3)
$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{t,i}+\Delta V^{\top}_{t,i})$ (4)
Here, $k=\min(m,n)$; ${\mathbf{u}}_{t}^{j}$ denotes the $j$-th column vector of matrix $U_{t}$, and ${\mathbf{v}}_{t}^{i}$ denotes the $i$-th row vector of matrix $V_{t}^{\top}$; $\Delta D_{t}\in(0,1)$ represents variation of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$; $U_{t}$ and $V_{t}^{\top}$ are decomposed from $W_{t}$ by performing SVD.
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer and without regularizer). See Section B.1 for training details of the analysis.
3.3 Analysis Result
This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect. The analysis data is illustrated in Figure 2.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: $\Delta D$ and $\Delta\Sigma$ of each trainable parameters during training steps. Numbers in the plot represent layer of the weight. Dots represent mean $\Delta D$ and $\Delta\Sigma$ at specific step.
The partial fine-tuning and LoRA methods analysis show that both methods exhibit significant adjustments in significant vectors $\Delta D$. This substantial alteration disrupts the characteristics of the pre-trained matrix and is likely to affect the model’s generalization ability. Moreover, the updates of all parameters in both partial fine-tuning and LoRA methods show a parallel updating pattern across weights in different layers, which emphasizes a restriction of these methods and may lead to a potentially disruptive modification of the generalization ability.
SORSA with regularizer shows $\Delta D$ and $\Delta\Sigma$ are in a much smaller range than other analysis methods. Moreover, different matrices demonstrated unrelated updating patterns, compared to all parallel-like other three methods, indicating that the updates in the SORSA are less constrained and potentially converge faster. In contrast, SORSA without the orthonormal regularizer exhibits a greater $\Delta D$ and $\Delta\Sigma$ alteration at each training step compared to SORSA with the regularizer. SORSA without regularizer also shows a linear-like updating pattern, similar to LoRA and partial FT. The result verifies the importance of regularizer in maintaining stability and minimizing the deviation from pre-trained matrices during the training process. These updating patterns indicate that SORSA maintains the characteristics of the pre-trained matrix better, thus potentially preserving the model’s generalization ability more effectively. This property allows SORSA adapters to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
4 SORSA: Singular Value and Orthonormal Regularized Singular Vector Adaptation
According to the definition of SVD in Section 3.1, given a rank $r$ where $r\ll k$, we could perform the low-rank approximation by selecting the first $r$ items on the diagonal of $\Sigma$, which is the first $r$ most significant singular values, and also select the first $r$ columns of $U$ and first $r$ rows of $V^{\top}$, which correspond to the selected singular values. By performing SVD low-rank approximation, we could get a low-rank matrix that preserves the largest significant values and vectors, containing the matrix’s most significant data.
Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principal weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
$W_{p}=U_{[:,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (5)
$W_{r}=U_{[:,r:]}\text{diag}(S_{[r:]})V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (6)
Here, $U$ represents the matrix of left singular vectors, $S$ represents the singular values, $\text{diag}(W)$ denotes a function to form a diagonal matrix from $W$, and $V$ represents the matrix of right singular vectors. We use PyTorch (Paszke et al., 2019) syntax to demonstrate matrix selection, where $[:,:r]$ denotes selecting the first $r$ columns of the matrix, and $[r:,:]$ denotes selecting the last $r$ rows of the matrix. We rewrite $U_{[:,:r]}$, $S_{[:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $S_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $S_{[:r]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $S_{r}$ and $V^{\top}_{r}$ correspondingly.
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\text{diag}(S_{p})$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\text{diag}(S_{p})^{\frac{1}{2}}$ and $B=\text{diag}(S_{p})^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $S_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
During training, $W_{r}$ remains frozen, and only $U_{p}$, $S_{p}$, and $V^{\top}_{p}$ are updated.
SORSA is defined as:
$\text{SORSA}(x):=x(W_{r}+W_{p})=xW_{r}+xU_{p}\text{diag}(S_{p})V^{\top}_{p}$ (7)
In our implementation, we use a optimized version of SORSA equation which results a much faster computation speed, and could approach to the speed of LoRA. See Appendix A for more details.
We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 5.
${\mathcal{L}}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V^{\top}_{p}V_{p}-I\|_{F}$ (8)
Where ${\mathcal{L}}_{reg}$ is the orthonormal regularizer loss, the $U_{p}$ and $V^{\top}_{p}$ are each orthonormal vectors in columns and rows, respectively, after initialization due to SVD’s property. The regularizer could keep their orthonormality during training.
Therefore, parameter updating of $W_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
$W_{p,t+1}=$ $W_{p,t}-\eta_{t}\nabla_{W_{p,t}}{\mathcal{L}}_{train}-\gamma_{t}\nabla_{W_{p,t}}{\mathcal{L}}_{reg}$ (9)
At training step $t$, $\nabla_{W_{p,t}}{\mathcal{L}}_{train}$ denotes the gradient of ${\mathcal{L}}_{train}$ respect to $W_{p,t}$, and $\nabla_{W_{p,t}}{\mathcal{L}}_{reg}$ denotes the gradient of ${\mathcal{L}}_{reg}$ respect to the orthonormal regularizer loss $W_{p,t}$. $\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss at step $t$, respectively.
We update the SORSA as Equation 10 for implementation simplicity.
$W_{p,t+1}=$ $W_{p,t}-\eta_{t}\left(\nabla_{W_{p,t}}{\mathcal{L}}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}{\mathcal{L}}_{reg}\right)$ (10)
$\eta_{d}$ is the maximum learning rate which is given from the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
5 Gradient Analysis
In this section, we present a comprehensive mathematical analysis of the SORSA method which mainly focus on the effect of orthonormal regularization. Our investigation elucidates the key optimization properties of SORSA, providing a theoretical foundation for its advantages. We explore four critical aspects: the convexity of the regularizer, the Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$, and the impact on the condition number of the optimization problem.
The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix C.
Our analysis reveals fundamental theoretical properties of SORSA, establishing its mathematical soundness and demonstrating its optimization advantages. We prove two key theorems that form the cornerstone of our theoretical framework:
Theorem 1.
The regularizer ${\mathcal{L}}_{reg}$ is convex.
Theorem 2.
The gradient of the regularizer ${\mathcal{L}}_{reg}$ is Lipschitz continuous.
These properties collectively ensure a well-behaved optimization landscape, facilitating stable and efficient training of SORSA adapters.
Building upon these foundational results, we further analyze the bounds of the hyperparameter $\gamma$, a critical factor in the performance of SORSA:
Theorem 3.
For convergence of gradient descent, the learning rate $\eta_{d}$ and regularization parameter $\gamma$ must satisfy:
$0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
This theorem provides crucial guidance for practitioners, offering a clear criterion for selecting appropriate values of $\gamma$ to ensure convergence of the gradient descent process.
To demonstrate the superior optimization properties of SORSA, we present a novel analysis of its condition number, a critical factor in determining convergence speed and stability. Our theoretical investigation reveals a significant improvement in the condition number compared to unregularized approaches, providing a mathematical foundation for SORSA’s enhanced performance.
We begin this analysis by establishing a key lemma that bounds the effect of the orthonormal regularizer on the singular values of the weight matrix:
Lemma 4.
Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\text{diag}(S_{p})^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\text{diag}(S_{p})^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer. For each singular value $\sigma_{i}$, the following bound holds:
$(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (12)
where $\epsilon$ is a small positive constant.
Lemma 4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
Theorem 5.
The orthonormal regularizer in SORSA can improve the condition number of the optimization problem over the course of training under certain conditions. Specifically:
At initialization ($t=0$):
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=1$ (13)
When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$. When $\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.
This theorem quantifies the improvement in the condition number achieved by SORSA, offering an explanation for its fast convergence. The proof leverages the effects of the orthonormal regularization to establish a tight bound on the condition number ratio. This theorem could also shows that training with the regularizer, the distribution of $W_{p}$ will be more evenly distributed due to a smaller ratio between $\sigma_{\max}(W_{p})$ and $\sigma_{\min}(W_{p})$, which means a better training stability.
Moreover, as mentioned in Büyükakyüz (2024), orthonormal matrices in neuron networks could improve gradient flow (Saxe et al., 2014; Arjovsky et al., 2016) and enhanced optimization landscape (Huang et al., 2018; Wisdom et al., 2016), which could also explain SORSA’s superior performance in convergence.
In conclusion, these theorems provide a mathematical foundation for the SORSA method. These theoretical guarantees validate the empirical success of SORSA but also provide valuable insights for future developments in PEFT methods.
6 Empirical Experiments
We conduct comparative experiments on different NLP tasks, including natural language generation (NLG) between SORSA, PiSSA (Meng et al., 2024), LoRA (Hu et al., 2021), and full parameter fine-tuning.
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023), RWKV6 7B (Peng et al., 2024), Mistral 7B v0.1 (Jiang et al., 2023) Gemma 7B (Gemma Team et al., 2024). We trained the models using first 100K data in MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We also trained the model on first 100K data in CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021). The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). All reported values are accuracy. See  Section B.2 for more details and hyperparameters of the training. We quoted some results of PiSSA, LoRA, and full parameter fine-tuning from Meng et al. (2024). For experiments, we conducted part of our experiments on a single NVIDIA A100-SXM4 (80GB) GPU, others on a single NVIDIA H100-SXM4 (80GB) GPU. See Table 1 for the results and Figure 3 for the loss and gradient norm comparison.
Model Method Trainable Parameters GSM-8K MATH HumanEval
Llama 2 7B Full FT 6738M 49.05 7.22 21.34
LoRA 320M 42.30 5.50 18.29
PiSSA 320M 53.07 7.44 21.95
SORSA 320M 56.03 10.36 24.39
RWKV6 7B LoRA 176M 8.04 7.38 15.24
PiSSA 176M 32.07 9.42 17.07
SORSA 176M 45.87 11.32 22.56
Mistral 7B Full FT 7242M 67.02 18.60 45.12
LoRA 168M 67.70 19.68 43.90
PiSSA 168M 72.86 21.54 46.95
SORSA 168M 73.09 21.86 47.56
Gemma 7B Full FT 8538M 71.34 22.74 46.95
LoRA 200M 74.90 31.28 53.66
PiSSA 200M 77.94 31.94 54.27
SORSA 200M 78.09 29.52 55.49
Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from Meng et al. (2024).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of RWKV6 7B and Llama 2 7B. LoRA and PiSSA curves of Llama 2 7B are from Meng et al. (2024).
The results showed that across all models tested, SORSA generally outperformed other methods. For mathematical evaluations, on Llama 2 7B, SORSA scored 56.03% on GSM-8K and 10.36% on MATH, significantly outperforming other methods; For the RWKV6 7B model, SORSA achieved 45.87% accuracy on GSM-8K and 11.32% on MATH, surpassing both PiSSA and LoRA; On Mistral 7B, SORSA reached 73.09% on GSM-8K and 21.86% on MATH, slightly edging out competitors; With Gemma 7B, while SORSA’s MATH score of 29.52% was lower than LoRA and PiSSA, it still led on GSM-8K with 78.09%. Moreover, on coding evaluations, SORSA consistently achieved the highest HumanEval scores across other tuning methods, demonstrating its effectiveness on diverse natural language tasks. Noticeably, RWKV6 7B using LoRA performed extremely below our expectation in GSM-8K test. We believe this may be due to our prompt for mathematical tests (See Section B.2 for more details) are inconsistent with the model’s original prompt, causing LoRA to fail in learning this prompting behavior.
The Figure 3 reveals that while SORSA and PiSSA exhibit nearly identical loss curves at the beginning, and even a little bit higher than PiSSA on RWKV-6 training. However, when training step approximately $t>300$, SORSA continues to maintain the decrease of its loss steadily. In contrast, LoRA and PiSSA shows a deceleration in its loss reduction. This supports Theorem 5, confirming that SORSA consistently outperforms both LoRA and PiSSA, especially at later stages of training.
However, due to the limitation of computing resources, we only trained and benchmarked a small number of tasks.
7 Conclusion
In this paper, we introduced SORSA, a novel parameter-efficient fine-tuning (PEFT) method designed to enhance the adaptation of large language models (LLMs) for downstream tasks. SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principal and residual components, only training the principal singular values and vectors while freezing the residuals. We implemented an orthonormal regularizer to maintain the orthonormality of singular vectors during training, ensuring efficient parameter updates and preserving the integrity of singular values.
Our experiments demonstrated that SORSA outperforms existing PEFT methods, such as LoRA and PiSSA, in both convergence speed and accuracy on the NLG tasks. Specifically, Llama 2 7B, tuned with SORSA, achieved significant improvements in the GSM-8K and MATH benchmarks, highlighting the effectiveness of our approach.
We adopted singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis.
Our gradient analysis provided a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape. This theoretical framework not only explains SORSA’s empirical superior performance but also offers valuable insights for future developments in adaptive learning algorithms.
SORSA retains the advantages of LoRA and variants, including low training VRAM requirements, no inference latency, and versatility across different neural network architectures. By offering a more efficient fine-tuning mechanism, SORSA presents a promising direction for future research and application in the field of LLMs.
Overall, SORSA gives a new perspective on parameter-efficient fine-tuning, showcasing exceptional efficiency and robust performance. It not only outperforms existing methods like LoRA and PiSSA in several downstream tasks but also maintains the practical benefits of low VRAM requirements, no inference latency, and ease of implementation. This innovative approach offers a promising direction for future research and practical applications in adapting pre-trained models, making it a pivotal development in the field.
8 Future Work
While SORSA demonstrates substantial improvements over existing PEFT methods, there are several avenues for future work to enhance its capabilities further and extend its applicability:
  • Comparison Between Different Rank: Find how the performance of SORSA varies with different ranks. This investigation would help determine the optimal rank for various tasks and model sizes, potentially revealing a trade-off between computational efficiency and model performance.
  • Extended Evaluation: Conduct broader evaluations of SORSA across different areas of benchmarks on models with more diverse sizes. This may able to provide a more comprehensive understanding of SORSA’s performance and versatility.
  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2020) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024) and QPiSSA (Meng et al., 2024). Quantization aims to reduce the memory footprint and computational requirements, making it feasible to deploy on devices with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, quantization of SORSA may able to efficiently provides powerful models, which could eventually facilitate AI for a broader range of practical applications.
By pursuing these directions, future work can build on the foundation laid by SORSA, pushing the boundaries of PEFT and enhancing the adaptability and performance of large language models across a wide range of applications. This will enable the development of more versatile downstream models, potentially expanding the impact of Machine Learning into more fields and becoming an integral part of people’s everyday lives.
9 Impact Statement
In this paper, we introduced an innovative PEFT method in the area of Machine Learning. Our approach significantly streamlined the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability. As we push the boundaries of what is possible with Machine Learning, it is essential to consider the broader impacts of these advancements on both the environment and ethical standards within the field.
9.1 Environmental Impact
SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B Dubey et al. (2024) model than full parameter fine-tuning. This significant reduction in hardware resource requirements also suggests a less energy consumption compared to full parameter fine-tuning methods. By enhancing efficiency, our approach could significantly contributes to reducing the carbon footprint of Machine Learning operations.
9.2 Ethical Concerns
The PEFT method, while efficient, raises critical ethical concerns regarding the security of built-in safety measures in AI models. As demonstrated in Lermen & Rogers-Smith (2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content. The ease and affordability of such methods underscore the vulnerability of safety protocols. It is imperative to develop robust safeguards that keep pace with technological advancements, ensuring that efficiency gains in model tuning do not compromise the ethical use of AI.
10 Acknowledgment
We would like to extend our gratitude to Prof. Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing the endorsement to the research.
Appendix A Faster SORSA Adapters
According to the definition of SORSA from Equation 7, because $\text{diag}(S_{p})$ is always a diagonal matrix, it is equivalent to:
$\text{SORSA}(x)=xW_{r}+x(U_{p}\odot S_{p})V_{p}^{\top}$ (14)
Where $\odot$ denotes element-wise multiplication, and $\text{diag}(W)$ denotes the diagonal elements of the matrix $W$.
This transformation allows us to reduce the computational complexity of SORSA adapters. In the original form, we have to perform twice matrix multiplications. However, in the Equation 14, we only have one matrix multiplication, and one element-wise multiplication. The time complexity of element-wise multiplication is $O(m^{2})$, which is much less than matrix multiplication’s complexity $O(m^{3})$. Therefore, while $\lim_{m\to\infty}$, the computation speed of SORSA adapters will be same as LoRA and PiSSA.
We performed a benchmark using PyTorch (Paszke et al., 2019) on a NVIDIA H100 SXM4 (80GB) GPU backed with CUDA, and Apple M2Pro CPU, in order to test the computation time between these two methods. See Figure 4 to see our results.
Refer to caption Refer to caption
Figure 4: Benchmark between two equations of SORSA
Appendix B Experiments Details
B.1 Analysis
For the singular values and vectors analysis in Section 3, we applied partial fine-tuning, LoRA and SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. We only calculated the loss on the response part. The models are trained with TF32 & BF16 (Wang & Kanwar, 2019) mix precision. See Table 2 for hyperparameters.
We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. In the analysis, LoRA and SORSA were only applied to q_proj and v_proj matrices, respectively. For Partial FT, we set model’s q_proj and v_proj matrices to trainable.
We also found we should only perform SVD for analysis using CPU, in order to get the precise analysis data.
Model Llama 2 7B
Method Partial FT LoRA SORSA (w/o reg) SORSA
Training
Mix-Precision TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16
Epoch 1 1 1 1
Batch Size 128 128 128 128
Max Length 512 512 512 512
Weight Decay 0 0 0 0
Warm-up Ratio 0.03 0.03 0.03 0.03
Learning Rate 2e-5 2e-5 2e-5 3e-5
Grad Clip False False False False
SORSA $\gamma$ N/A N/A 0 5e-4
Rank N/A 128 128 128
Table 2: Hyperparameters for the analysis
B.2 NLG Experiments
For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023), RWKV6 7B (Peng et al., 2024), Mistral 7B v0.1 (Jiang et al., 2023) Gemma 7B (Gemma Team et al., 2024). models by SORSA. For GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) evaluations, we trained those models with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. For HumanEval (Chen et al., 2021) evaluation, we use the first 100K data in CodeFeedback Filtered Instruction(Zheng et al., 2024) dataset.
We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA adapters were applied on all linear matrices in every layer. We only calculated the loss on the response part. The models are loaded in FP32, and trained with TF32 & BF16 mix precision. In our experiments, we selected a higher learning rate for SORSA compared to other methods to counterbalance the negative effect of orthonormal regularizer on optimizing toward lower training loss. See Table 3 for hyperparameters. See LABEL:lst:math for the prompt we used in GSM-8K and MATH evaluations, and LABEL:lst:humaneval for the prompt we used for HumanEval tests.
Model Llama 2 7B Mistral 7B Gemma 7B RWKV6 7B RWKV6 7B
Method SORSA SORSA SORSA SORSA LoRA PiSSA
Training
Mix-Precision TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16
Epoch 1 1 1 1 1
Batch Size 128 128 128 128 128
Max Length 512 512 512 512 512
Weight Decay 0 0 0 0 0
Warm-up Ratio 0.03 0.03 0.03 0.03 0.03
Learning Rate 3e-5 3e-5 3e-5 3e-5 2e-5
Grad Clip 1.0 1.0 1.0 1.0 1.0
SORSA $\gamma$ 4e-4 4e-4 4e-4 4e-4 N/A
Rank 128 64 64 64 64
Evaluating
Precision BF16 BF16 BF16 FP32 FP32
Sampling False
Top-P 1.0
Max Length GSM-8K: 1024 MATH: 2048 HumanEval: 2048
Table 3: Hyperparameters of training on models for GSM-8K and MATH
1Below is an instruction that describes a task. Write a response that appropriately completes the request.
2
3### Instruction:
4{question}
5
6### Response: Lets think step by step.
Listing 1: Prompt used for GSM-8K and MATH.
1@@ Instruction
2Here is the given code to do completion:
3‘‘‘python
4{question}
5‘‘‘
6
7Please continue to complete the function with python programming language. You are not allowed to modify the given code and do the completion only.
8
9Please return all completed codes in one code block.
10This code block should be in the following format:
11’’’python
12# Your codes here
13’’’
14
15@@ Response
Listing 2: Prompt used for HumanEval evaluation.
Appendix C Proofs
See 1
Proof.
We prove this in two steps:
First, we show that $f(U_{p})=\|U^{\top}_{p}U_{p}-I\|_{F}$ is convex. Then, we prove that $g(V_{p})=\|V_{p}V^{\top}_{p}-I\|_{F}$ is convex.
Since the sum of convex functions is convex, this will establish the convexity of ${\mathcal{L}}_{reg}$.
Let $U_{p},W\in\mathbb{R}^{m\times r}$. The Hessian of $f$ at $U_{p}$ in the direction $W$ is given by:
$\begin{split}\nabla^{2}f(U_{p})[W,W]&=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(f(U_{p}+\epsilon W)-2f(U_{p})+f(U_{p}-\epsilon W)\Big)\\ &=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(\|(U_{p}+\epsilon W)^{\top}(U_{p}+\epsilon W)-I\|_{F}\\ &\quad-2\|U^{\top}_{p}U_{p}-I\|_{F}+\|(U_{p}-\epsilon W)^{\top}(U_{p}-\epsilon W)-I\|_{F}\Big)\\ &=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(\|U^{\top}_{p}U_{p}+\epsilon(U^{\top}_{p}W+W^{\top}U_{p})+\epsilon^{2}W^{\top}W-I\|_{F}\\ &\quad-2\|U^{\top}_{p}U_{p}-I\|_{F}+\|U^{\top}_{p}U_{p}-\epsilon(U^{\top}_{p}W+W^{\top}U_{p})+\epsilon^{2}W^{\top}W-I\|_{F}\Big)\\ &=2\|W^{\top}W\|_{F}\end{split}$ (15)
Since $\|W^{\top}W\|_{F}\geq 0$ for all $W$, we have $\nabla^{2}f(U_{p})[W,W]\geq 0$, which proves that $f$ is convex.
The proof for $g(V_{p})$ follows the same steps as for $f(U_{p})$, leading to the same conclusion.
Therefore, both $f(U_{p})$ and $g(V_{p})$ are convex, and consequently, ${\mathcal{L}}_{reg}$ is convex. ∎
See 2
Proof.
To prove Lipschitz continuity, we need to show that there exists a constant $L>0$ such that for any two pairs of matrices $(U_{1},V_{1})$ and $(U_{2},V_{2})$:
$|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|\leq L(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})$ (16)
Let’s break this down into two parts:
First, consider $|\|U_{1}^{\top}U_{1}-I\|_{F}-\|U_{2}^{\top}U_{2}-I\|_{F}|$:
$\begin{split}|\|U_{1}^{\top}U_{1}-I\|_{F}-\|U_{2}^{\top}U_{2}-I\|_{F}|&\leq\|(U_{1}^{\top}U_{1}-I)-(U_{2}^{\top}U_{2}-I)\|_{F}\\ &=\|U_{1}^{\top}U_{1}-U_{2}^{\top}U_{2}\|_{F}\\ &=\|U_{1}^{\top}U_{1}-U_{1}^{\top}U_{2}+U_{1}^{\top}U_{2}-U_{2}^{\top}U_{2}\|_{F}\\ &\leq\|U_{1}^{\top}(U_{1}-U_{2})\|_{F}+\|(U_{1}^{\top}-U_{2}^{\top})U_{2}\|_{F}\\ &\leq\|U_{1}^{\top}\|_{F}\|U_{1}-U_{2}\|_{F}+\|U_{1}-U_{2}\|_{F}\|U_{2}\|_{F}\\ &\leq(\|U_{1}\|_{F}+\|U_{2}\|_{F})\|U_{1}-U_{2}\|_{F}\end{split}$ (17)
Here, we’ve used the triangle inequality and the sub-multiplicative property of the Frobenius norm.
Similarly for $V$:
$|\|V_{1}^{\top}V_{1}-I\|_{F}-\|V_{2}^{\top}V_{2}-I\|_{F}|\leq(\|V_{1}\|_{F}+\|V_{2}\|_{F})\|V_{1}-V_{2}\|_{F}$ (18)
Combining these results:
$\begin{split}|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|&\leq(\|U_{1}\|_{F}+\|U_{2}\|_{F})\|U_{1}-U_{2}\|_{F}\\ &\quad+(\|V_{1}\|_{F}+\|V_{2}\|_{F})\|V_{1}-V_{2}\|_{F}\\ &\leq\max(\|U_{1}\|_{F}+\|U_{2}\|_{F},\|V_{1}\|_{F}+\|V_{2}\|_{F})\\ &\quad\cdot(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})\end{split}$ (19)
Let $L_{reg}=\max(\|U_{1}\|_{F}+\|U_{2}\|_{F},\|V_{1}\|_{F}+\|V_{2}\|_{F})$. This $L$ is finite because the Frobenius norms of $U$ and $V$ are bounded (they represent orthonormal matrices in the ideal case).
Therefore, we have shown that:
$|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|\leq L_{reg}(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})$ (20)
This proves that ${\mathcal{L}}_{reg}$ is Lipschitz continuous with Lipschitz constant $L_{reg}$. ∎
See 3
Proof.
For gradient descent to converge while the $\eta_{t}=\eta_{d}$, we need:
$\eta_{d}<\frac{2}{L}=\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (21)
Rearranging this inequality:
$\begin{split}\eta_{d}(L_{train}+\frac{\gamma}{\eta_{d}}L_{reg})&<2\\ \eta_{d}L_{train}+\gamma L_{reg}&<2\\ \gamma L_{reg}&<2-\eta_{d}L_{train}\\ \gamma&<\frac{2-\eta_{d}L_{train}}{L_{reg}}\end{split}$ (22)
We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$. Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$. Therefore, we could assume $L_{reg}\propto\eta_{d}$. Moreover, the $\gamma$ must not be negative or the regularization term would impact negatively on the purposes it is supposed to have. Therefore, we can rewrite the inequality as:
$0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (23)
This establishes the upper bound for $\gamma$. ∎
See 4
Proof.
First, let’s consider the effect of the orthonormal regularizer. The regularizer aims to make $U_{p}^{\top}U_{p}\approx I$ and $V_{p}V_{p}^{\top}\approx I$. We can quantify this approximation as:
$\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}\leq\epsilon_{\nabla}$ (24)
where $\epsilon_{\nabla}>0$ is a small constant.
Now, let’s consider the difference between $W_{p}^{\text{unreg}}$ and $W_{p}^{\text{reg}}$:
$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}=\gamma\epsilon_{\nabla}$ (25)
Now, we can use Weyl’s inequality (Weyl, 1912), which states that for matrices A and B:
$|\sigma_{i}(A+B)-\sigma_{i}(A)|\leq\|B\|_{2}\leq\|B\|_{F}$ (26)
Applying this to our case, with $A=W_{p}^{\text{unreg}}$ and $B=W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}$:
$|\sigma_{i}^{\text{reg}}-\sigma_{i}^{\text{unreg}}|\leq\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}\leq\gamma\epsilon_{\nabla}$ (27)
Let $\epsilon=\gamma\epsilon_{\nabla}$. Then we have:
$-\epsilon\leq\sigma_{i}^{\text{reg}}-\sigma_{i}^{\text{unreg}}\leq\epsilon$ (28)
Rearranging this inequality gives us our desired bound:
$(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (29)
See 5
Proof.
Let $W_{p}=U_{p}\text{diag}(S_{p})V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
$\kappa(W_{p})=\frac{\sigma_{\max}(W_{p})}{\sigma_{\min}(W_{p})}$ (30)
where $\sigma_{\max}$ and $\sigma_{\min}$ are the maximum and minimum singular values of $W_{p}$.
At initialization ($t=0$): Due to SVD initialization, $U_{p}$ and $V_{p}^{\top}$ are perfectly orthonormal, so:
$\kappa(U_{p}^{\text{unreg}})=\kappa((V_{p}^{\text{unreg}})^{\top})=\kappa(U_{p}^{\text{reg}})=\kappa((V_{p}^{\text{reg}})^{\top})=1$ (31)
And $\epsilon_{0}=0$, $\delta_{1,0}=\delta_{2,0}=0$. Therefore:
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\text{diag}(S_{p})^{\text{reg}})}{\kappa(\text{diag}(S_{p})^{\text{unreg}})}=1$ (32)
During training ($t>0$): As training progresses, $U_{p}$ and $V_{p}^{\top}$ deviate from orthonormality in the unregularized case. We quantify this deviation:
$\|U_{p}^{\top}U_{p}-I\|_{F}\leq\epsilon_{1,t}$ (33)
$\|V_{p}^{\top}V_{p}-I\|_{F}\leq\epsilon_{2,t}$ (34)
where $\epsilon_{1,t},\epsilon_{2,t}>0$ increase over time t.
For the regularized matrices, we can bound their condition numbers:
$\kappa(U_{p}^{\text{reg}})\leq 1+\delta_{1,t}$ (35)
$\kappa(V_{p}^{\text{reg}})\leq 1+\delta_{2,t}$ (36)
where $\delta_{1,t},\delta_{2,t}$ are small positive numbers that remain bounded due to the regularization.
From the Lemma 4, we arrive at:
$\begin{split}\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1-\epsilon_{t}}{1+\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\leq\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}\leq\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\end{split}$ (37)
As training continues, in the unregularized case, $\kappa(U_{p,\text{unreg}})$ and $\kappa(V_{p,\text{unreg}}^{\top})$ tend to increase as $U_{p}$ and $V_{p}^{\top}$ deviate further from orthonormality. On the other hand, $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will approach to 1 because of the reinforcement in orthonormality. Therefore, $\exists c>0$, while $t>c$:
$\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})>(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ (38)
will hold true.
Therefore, while $t>c$, we have:
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}<1$ (39)
indicating an improvement in the condition number. ∎